At 9am a plane leaves chicago heading west. At 11 am a second plane leaves chicago heading west on the same flight path and treveling 50mph faster than the first plane. the second plane overtakes the first plane in 5 hours. What is the rate of the first plane?
Answer by checkley77(12844)

7R=5(R+50)
7R=5R+250
7R-5R=250
2R=250
R=250/2
R=125 MPH FOR THE FIRST PLANE.
125=50=175 MPH FOR THE SECOND PLANE.
PROOF:
7&125=5(125+50)
875=5*175
875=875 IS THE DISTANCE BOTH PLANES COVERED.